Fermi National Laboratory


Accelerator Update for November 16, 2005 - November 18, 2005

Wednesday November 16

The day shift began with interlock technicians conducting safety system checks of the accelerators, with Operations monitoring a small Antiproton Source (Pbar) stack, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) waiting to access their collision halls, with MiniBooNE, NuMI, and the Meson MCenter and MTest experiments down due to safety system checks.

Operations began preparing CDF and D0 for an access at 8:23 AM.

The Booster safety testing ended at 8:55 AM.

A Linac expert reported at 9:19 AM that they were beginning to repair a RF station (KRF6) LCW leak.

At 9:24 AM, LRF5 tripped off and wouldn’t reset. Linac techs investigated.

Cryo system techs began wet engine maintenance around the Tevatron (TeV) at 9:33 AM.

The Pbar pre-target safety system testing ended at 11:06 AM.

The MI safety system testing ended at 1:46 PM.

At 2:23 PM, Operations noticed that a vacuum pump had tripped off in the TeV near the sector B1 separators. Vacuum techs will access the area after the safety system tests end and check for a vacuum leak.

The TeV safety system testing ended at 3:58 PM.

Operations reported at 4:28 PM that Linac techs had completed repairs on LRF5. Linac resumed sending beam.

Operations established beam to NuMI at 4:31 PM.

Pbar resumed stacking at 4:39 PM.

At 5:10 PM, LRF5 tripped off. It reset okay.

Operations sent protons to the Recycler at 5:39 PM.

Vacuum techs reported at 9:17 PM that they found a B11 separator vacuum leak.

Operations prepared the Recycler to stash antiprotons at 10:42 PM. The stash ended at 10:57 PM after six transfers. Pbar resumed stacking at 11:11 PM.

Thursday November 17

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE and NuMI taking beam.

Operations began the TeV turn on process at 1:13 AM. They discovered a ground fault indication on a sector D0 quadrupole (Q2-4). Operators called in experts.

The Recycler began preparing to stash around 3:08 AM. The stash ended. Pbar resumed stacking at 4:10 AM. MCenter and MTest resumed taking beam.

At 7:26 AM, while a Switchyard (SY) expert was working at increasing MTest’s intensity, their intensity dropped. Operators investigated and discovered that vacuum valves had closed in the P3 transfer line.

Power supply experts reported at 8:10 AM that they had completed repairs on the D0 Q2-4 ground fault.

Operations began another TeV turn on at 8:27 AM.

The Recycler began preparing to stash around 8:32 AM. The stash ended. Pbar resumed stacking at 9:55 AM.

A TeV expert began a thirty-minute Tevatron Electron Lens study at 10:06 AM.

MCenter and MTest resumed taking beam at 10:18 AM.

At 10:35 AM, a TeV quadrupole power supply (C:BQ7) tripped off on an input power fault. Operations contacted experts.

At 11:45 AM, the P1 transfer line suffered a vacuum burst. Operators restarted a vacuum pump.

Power supply experts reported at 12:01 PM that they had repaired C:BQ7.

Operations put the TeV through a dry squeeze at 12:02 PM.

Pbar studies ended at 12:53 PM.

The Recycler began preparing to stash around 12:54 PM.

At 1:12 PM, the P3 transfer line vacuum valves closed. Operations saw no reason for this, so they contacted experts. This holds off SY.

At 1:16 PM, a TeV quadrupole power supply (C:B0QT3) tripped off. An operator contacted a power supply expert.

The stash ended at 1:37 PM.

Pbar resumed stacking at 1:54 PM.

The P3 vacuum work ended and MCenter and MTest resumed taking beam around 2:36 PM.

Operations began preparing the TeV for an access at 2:59 PM. Vacuum techs will leak-check the B1 separator again.

D0 accessed their collision hall at 3:24 PM and came out about an hour and a half later.

At 4:54 PM, a vacuum burst in Booster caused BRF7 & 8 to trip off. Everything reset okay.

Operations put the TeV through a dry squeeze at 9:04 PM.

Operations put the TeV through a wet squeeze at 10 PM.

Operations put the TeV into shot setup at 11:36 PM.

Friday November 18

The midnight shift began with the TeV in shot setup, and with MiniBooNE and NuMI taking beam.

At 12:02 PM, Operations reported a power glitch at the Master Substation. Operations contacted the Duty Electrician.

At 12:29 AM, the P150 transfer line suffered a vacuum burst and lost its vacuum beam permit. The vacuum recovered within two minutes.

At 1:30 AM, after the first transfer of antiprotons the Recycler RF structure went to a mixed state. Operations saw no way of recovering from this so they called in an expert. This put the TeV shot setup on hold.

The Duty Electrician reported at 2:05 AM that the power glitch was caused by a feeder (#37) fault. This fault left the Tag Photon Lab, the Magnet Storage Building, the Proton Assembly Building, and Wideband without power. Later, generators were installed to supply power to these areas.

Operations put the TeV through a dry squeeze at 2:09 AM.

At 2:32 AM, BRF16 tripped off due to sparking. The station continued to spark so operators bypassed the station.

MTest resumed taking beam around 3:54 AM. MCenter was still off due to its MS6 LCW system being down, which was fed from feeder 37.

Operations put the TeV back into shot setup at 4:12 AM.

The MS6 LCW system came back on at 5:34 AM.

Operations prepared MI at 6:32 AM to access MI-40 to check its sump pump. They found that the high-level float assembly had slipped down.

Operations established store 4506 at 6:53 AM with an initial luminosity of 64.13E30.

At 7:50 AM, a water tech found a LCW leak on a MI quadrupole (MI-212). They began repairs.

The Plans for Friday and the Weekend

The plans for today are to stash, stack, and store. The Run Coordinator will allow a Pbar and TeV study today. The present store will be terminated and a new store established later this evening.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 11/18/2005   email Fermilab